Find the measure of two supplementary angles if the measure of the larger angle is 20 more than the smaller angle. I have tried the way I would try it but the larger angle came out equaling less and I have no clue how to find the answer
Answer by stanbon(75887)

Find the measure of two supplementary angles if the measure of the larger angle is 20 more than the smaller angle
---
1st: x
2nd: 180-x
----
Note: At this point you do not know which angle is larger.
----
Equation:
180-x = x +20
-----
Note: Now you are saying 180-x is 20 larger than "x".
-----
2x = 160
1st: x = 80
2nd: 180-x = 100
------------------
Cheers,
Stan H.